If displacement of particle is x=sin(5pi t) +4cos(5pi t). Then amplitude is

Grade:12

3 Answers

Sandeep Bhamoo
5 years ago
Amplitude is the maximum value of displacement (here given by x). For maximum value of x , (dx)/dt = 0 On solving we get time at which x is max. it is tan(5πt)=4 on putting value of t , x = 8/(√17). Please do approve my answer if correct.
Khimraj
5 years ago
Maximum and minimum value of x is ^{\sqrt{17}} and -^{\sqrt{17}}. So particle is oscillating between these positions. So amplitude will be given by
2A = ^{\sqrt{17}} - (-^{\sqrt{17}}) = 2^{\sqrt{17}}.
So A = ^{\sqrt{17}}.
Hope it clears.
2A = 4 - (-3) = 
Khimraj
5 years ago
Last line printed by mistake...........................................................................